# PlayerOne: Egocentric World Simulator

### v1

@@ 3.2 Model Components
$[\mathbf{u}]_{\times}=\left[\begin{array}[]{ccc}0&-u_{z}&u_{y}\\ u_{z}&0&-u_{x}\\ -u_{y}&u_{x}&0\end{array}\right]$ (4)
@@ 4.1 Experimental Setting
Implementation details. We choose Wanx2.1 1.3B wan2025 as the base generator. We set the LoRA rank and the update weight of the matrices as 128 and 4 respectively and initialize its weight following  wan2025 . The inference step and the learning rate are set as 50 and $1\times 10^{-5}$ respectively, where the Adam optimizer and mixed-precision bf16 are adopted. The cfg of 7.5 is used. We train our model for 100,000 steps on 8 NVIDIA A100 GPUs with a batch size of 56 and sample resolution of 480$\times$480. The generated video runs at eight frames per second, and we utilize 49 video frames (6 seconds) for training. After distillation, our method can achieve 8 FPS to generate the desired results. All the action videos in this paper are shot with the front camera.
@@
Metrics. On our constructed benchmark, for evaluation of alignment with the given text descriptions, we calculate both CLIP-Score and DINO-Score, where PSNR, SSIM, and LPIPS zhang2018unreasonableeffectivenessdeepfeatures are employed to evaluate the video fidelity of the generated video. Besides, we calculate the frame consistency to evaluate the temporal coherence and consistency of the generated video frames over time. We further utilize a 3D hand pose estimation model prakash20243dhandposeestimation to estimate the hand pose of the generated videos and use the results of the ground truth video as the labels. Afterward, we follow  prakash20243dhandposeestimation to calculate two metrics: (1) Mean Per-Joint Position Error (MPJPE): the L2 distance between the predicted and ground truth joints for each hand after subtracting the root joint. (2) Mean Relative-Root Position Error (MRRPE): the metric distance between the root joints of the left hand and right hand.
@@ 4.2 Ablation Study
DINO-Score ($\uparrow$) CLIP-Score ($\uparrow$) MPJPE ($\downarrow$) MRRPE ($\downarrow$) PSNR($\uparrow$) FVD ($\downarrow$) LPIPS($\downarrow$)
Baseline 51.3 65.6 376.14 341.01 35.6 394.16 0.1421
+ Pretrain 56.6 74.4 258.05 232.17 41.2 301.32 0.1146
+ Pretrain&ControlNet 57.1 75.2 241.73 218.46 42.8 287.52 0.1103
+ Pretrain&Entangled 58.0 76.3 235.12 212.53 43.9 279.41 0.1060
+ Pretrain&PMI (No Camera) 60.7 79.8 183.25 196.35 45.6 257.04 0.0902
+ Pretrain&PMI 62.5 81.3 156.76 175.18 48.3 245.72 0.0839
+ Pretrain&PMI&Filtering 64.2 83.8 141.56 163.04 49.1 230.50 0.0782
+ Pretrain&PMI&Filtering&Recon(No Adapter) 62.7 81.6 176.23 180.10 47.3 240.17 0.0919
+ Pretrain&PMI&Filtering&Recon(DUSt3R) 67.5 87.7 129.08 152.22 52.2 228.20 0.0685
@@ 4.3 Comparison with State-of-the-arts
DINO-Score ($\uparrow$) CLIP-Score ($\uparrow$) MPJPE ($\downarrow$) MRRPE ($\downarrow$) PSNR($\uparrow$) FVD ($\downarrow$) LPIPS($\downarrow$)
Aether aether 38.0 64.2 415.70 431.05 38.1 397.40 0.1856
Cosmos(Diff-7B) nvidia2025cosmosworldfoundationmodel 45.3 70.3 301.92 324.12 43.7 346.09 0.1630
Cosmos(Diff-14B) nvidia2025cosmosworldfoundationmodel 51.6 79.7 256.73 253.06 47.8 302.17 0.1351
PlayerOne(ours) 67.8 88.2 127.16 151.62 52.6 226.12 0.0663

### v2

@@ 3.2 Model Components
$\mathbf{u}]_{\times}=\left[\begin{array}[]{ccc}0&-u_{z}&u_{y}\\ u_{z}&0&-u_{x}\\ -u_{y}&u_{x}&0\end{array}\right]$ (4)
@@ 4.1 Experimental Setting
Implementation details. We choose Wanx2.1 1.3B wan2025 as the base generator. We set the LoRA rank and the update weight of the matrices as 128 and 4 respectively and initialize its weight following  wan2025. The inference step and the learning rate are set as 50 and $1\times 10^{-5}$ respectively, where the Adam optimizer and mixed-precision bf16 are adopted. The cfg of 7.5 is used. We train our model for 100,000 steps on 8 NVIDIA A100 GPUs with a batch size of 56 and sample resolution of $\times480$. The generated video runs at eight frames per second, and we utilize 49 video frames (6 seconds) for training. After distillation, our method can achieve 8 FPS to generate the desired results. All the action videos in this paper are shot with the front camera.
@@
Metrics. On our constructed benchmark, for evaluation of alignment with the given text descriptions, we calculate both CLIP-Score and DINO-Score, where SSIM, and LPIPS zhang2018unreasonableeffectivenessdeepfeatures are employed to evaluate the video fidelity of the generated video. Besides, we calculate the frame consistency to evaluate the temporal coherence and consistency of the generated video frames over time. We further utilize a 3D hand pose estimation model prakash20243dhandposeestimation to estimate the hand pose of the generated videos and use the results of the ground truth video as the labels. Afterward, we follow  prakash20243dhandposeestimation to calculate two metrics: (1) Mean Per-Joint Position Error (MPJPE): the L2 distance between the predicted and ground truth joints for each hand after subtracting the root joint. (2) Mean Relative-Root Position Error (MRRPE): the metric distance between the root joints of the left hand and right hand.
@@ 4.2 Ablation Study
DINO-Score ($\uparrow$) CLIP-Score ($\uparrow$) MPJPE ($\downarrow$) MRRPE ($\downarrow$) FVD ($\downarrow$) LPIPS($\downarrow$)
Baseline 51.3 65.6 376.14 341.01 394.16 0.1421
+ Pretrain 56.6 74.4 258.05 232.17 301.32 0.1146
+ Pretrain&ControlNet 57.1 75.2 241.73 218.46 287.52 0.1103
+ Pretrain&Entangled 58.0 76.3 235.12 212.53 279.41 0.1060
+ Pretrain&PMI (No Camera) 60.7 79.8 183.25 196.35 257.04 0.0902
+ Pretrain&PMI 62.5 81.3 156.76 175.18 245.72 0.0839
+ Pretrain&PMI&Filtering 64.2 83.8 141.56 163.04 230.50 0.0782
+ Pretrain&PMI&Filtering&Recon(No Adapter) 62.7 81.6 176.23 180.10 240.17 0.0919
+ Pretrain&PMI&Filtering&Recon(DUSt3R) 67.5 87.7 129.08 152.22 228.20 0.0685
@@ 4.3 Comparison with State-of-the-arts
DINO-Score ($\uparrow$) CLIP-Score ($\uparrow$) MPJPE ($\downarrow$) MRRPE ($\downarrow$) FVD ($\downarrow$) LPIPS($\downarrow$)
Aether aether 38.0 64.2 415.70 431.05 397.40 0.1856
Cosmos(Diff-7B) nvidia2025cosmosworldfoundationmodel 45.3 70.3 301.92 324.12 346.09 0.1630
Cosmos(Diff-14B) nvidia2025cosmosworldfoundationmodel 51.6 79.7 256.73 253.06 302.17 0.1351
PlayerOne(ours) 67.8 88.2 127.16 151.62 226.12 0.0663